OpenStudy (anonymous):

ok so im getting this wrong the derivitive of ln 5x = to what? 1/5x? and the derivitve of Sqrt 5x = (5x)^1/2= 1/2(5x)^-1/2?? not sure how or where i went wrong

OpenStudy (anonymous):

is that even right?

OpenStudy (anonymous):

d/dx ln(5x) = (1/x)

OpenStudy (anonymous):

so 5 is gone?

OpenStudy (anonymous):

you nearly got the sqrt(5x) correct, you are mising a factor of 5 from the chain rule

OpenStudy (anonymous):

ok so how i do that?

OpenStudy (anonymous):

\[\frac{d}{dx} \ln(f(x)) = \frac{ f ' (x) }{f(x) }\]

OpenStudy (anonymous):

\[\frac{d}{dx} \ln(5x) = \frac{5}{5x} = \frac{1}{x}\]

OpenStudy (anonymous):

o ok...what about the sqrt(5x)..

OpenStudy (anonymous):

y= (5x)^(1/2) apply chain rule bring power down in front, leave the inside of bracket alone , reduce the power by 1 , then multiply the derivative of the inside of the bracket.

OpenStudy (amistre64):

Dx(ln(x)) = |Dx/x|

OpenStudy (anonymous):

thats what i did to become this: \[1/2(5x)^{-1/2}\]

OpenStudy (anonymous):

no

OpenStudy (anonymous):

is that righht?

OpenStudy (anonymous):

did u multiply by derivativbe of the inside?

OpenStudy (anonymous):

u are missing a factor of 5

OpenStudy (anonymous):

i dont think i was taught that yet...

OpenStudy (anonymous):

what you mean factors of 5

OpenStudy (anonymous):

just means multiply it by 5

OpenStudy (amistre64):

Since 5x plays an important part of the equation; it has to be accounted for in the derivative as well; by multiplying it into the derivative thru the "chain rule"

OpenStudy (anonymous):

but lim x approaches infinity.....

OpenStudy (anonymous):

can you show so i get the idea...

OpenStudy (amistre64):

Dx[F(g(x))] = F'(g(x)) * g'(x)

OpenStudy (amistre64):

F(x) = sqrt(x) g(x) = 5x Dx[sqrt(5x)] = 1/(2sqrt(5x)) * 5 = 5/2sqrt(5x)

OpenStudy (anonymous):

alternatively let u= 5x , then \[\frac{dy}{dx} = \frac{dy}{du} \times \frac{du}{dx}\]

OpenStudy (anonymous):

its easier to do it that way for beginners

OpenStudy (amistre64):

both explanations have their pros and cons :)

OpenStudy (anonymous):

so its 25x?

OpenStudy (anonymous):

after a while you can just do it

OpenStudy (anonymous):

no

OpenStudy (anonymous):

lol wow....i am totally lost

OpenStudy (amistre64):

\[\sqrt{5x}\implies\frac{5}{2\sqrt{5x}}\]

OpenStudy (anonymous):

I never go back to the u substitutions , I just remeber the chant "bring power down in front, leave the inside of bracket alone , reduce the power by 1 , then multiply the derivative of the inside of the bracket. "

OpenStudy (anonymous):

its too long, a waste of time in an exam

OpenStudy (amistre64):

its like peeling away the layers of an onion \[[F(g(h(x)))]' = F'(g(h(x)))*g'(h(x))*h'(x)*x'\]

OpenStudy (anonymous):

what the....? never even saw that.....

OpenStudy (anonymous):

mmm now i remeber!!